### v1

@@ Abstract
A graph $H$ is an induced minor of a graph $G$ if $H$ can be obtained from $G$ by a sequence of edge contractions and vertex deletions. Otherwise, $G$ is $H$-induced minor-free. In this paper, we prove that $K_{2,3}$-induced minor-free graphs admit a quasi-isometry with additive distortion to graphs with tree-width at most two. Our result implies that a recent conjecture of Nguyen et al. [Coarse tree-width (2025)] holds for $K_{2,3}$-induced minor-free graphs.
@@ 1 Introduction
When is a graph quasi-isometric (Definition 1) to a much simpler host graph? This question is central to coarse graph theory [19]. When the host graphs are restricted to the class of graphs with bounded tree-width, remarkably, the answer is provided by Nguyen, Scott & Seymour [22] and independently by Hickingbotham [21]. Specifically, they proved the following:
@@
When quasi-isometric maps with only additive distortions (Definition 2) are allowed, the situation is much less understood. Brandstädt et al. [5] proved that chordal graphs (i.e., graphs without any induced cycle of length greater than three) admit a quasi-isometry to trees with additive distortion at most $2$. This result has been later generalized to $k$-chordal graphs [10], and more generally to graphs of bounded treelength [16]. Berger & Seymour [3] characterized graphs admitting quasi-isometry with additive distortion to trees. See [17] for a comprehensive overview of graphs admitting a quasi-isometry with additive distortion to trees.
@@
In this paper, we consider the class of $K_{2,3}$-induced minor-free graphs. A graph $H$ is an induced minor of a graph $G$ if $H$ can be obtained from $G$ by a sequence of edge contractions and vertex deletions. Otherwise, $G$ is $H$-induced minor-free. The class of $K_{2,3}$-induced minor-free graphs have been studied extensively in structural graph theory [14, 15]. $K_{2,3}$-induced minor-free graphs generalize many popular graph classes. Examples include outerplanar graphs, circular-arc graphs [18], chordal graphs, universally signable graph [13] etc. Coarse generalizations of $K_{2,3}$-induced minor-free graphs have been studied [11]. A recent result of Albrechtsen et al. [2] implies that $K_{2,3}$-induced minor-free graphs are quasi-isometric to graphs with tree-width at most two. We show that 2 holds for $K_{2,3}$-induced minor-free graphs. Specifically, we prove the following theorem.
@@
There is a constant $c$ such that, for each $K_{2,3}$-induced minor-free graph $G$, there is a graph $H$ with $V(G)=V(H)$, tree-width at most two, and $\left|d_{G}\left(u,v\right)-d_{H}\left(u,v\right)\right|\leq c$ for any $u,v\in V(G)$.
@@
In the above theorem, $d_{G}\left(.,.\right)$ and $d_{H}\left(.,.\right)$ denote the standard graph metric in $G$ and $H$, respectively. Below, we provide an informal description of the tools used in the proof of Theorem 3. Then, we provide an overview of the proof of Theorem 3.
@@
Important tools: The proof of Theorem 3 uses layering partition [5, 10] and the notion of strong isometric path complexity [6] of graphs. Layering partitions have been used to solve numerous problems in metric graph theory [5, 10, 11, 17]. See Section 2.2 for definitions. Informally, in a layering partition, the vertex set of a graph is partitioned into clusters w.r.t a fixed root vertex. Two vertices are in the same cluster if they are at a distance $k$ from the root for some $k\geq 0$ and are end-vertices of a path that is disjoint from the ball of radius $k-1$ centered at the root vertex. It was shown in [5, 10] that the adjacency relation between the clusters is captured by a tree, known as the layering tree. Moreover, for a fixed root vertex, the corresponding layering tree can be constructed in linear time [10].
@@
The strong isometric path complexity is a recently introduced graph invariant that captures how arbitrary isometric paths of a graph can be viewed as a union of few “rooted" isometric paths (i.e. isometric paths with a common end-vertex). See Definition 3. Informally, if the strong isometric path complexity of a graph $G$ is at most $k$, then for any isometric path $P$ of $G$ and any vertex of $v\in V(G)$, the vertices of $P$ can be covered with $k$-many $v$-rooted isometric paths of $G$. Originally, strong isometric path complexity and its weaker variant isometric path complexity were introduced to propose approximation algorithms for the problem Isometric Path Cover, where the objective is to find a minimum-cardinality set of isometric paths so that each vertex of the graph belongs to at least one of the solution paths. Isometric Path Cover has been studied recently in the algorithmic graph theory community [7, 6, 9]. It was shown that Isometric Path Cover admit a polynomial-time $c$-approximation algorithm on graphs with strong isometric path complexity at most $c$ [6] (while being NP-hard even on chordal graphs whose strong isometric path complexity is at most four [7, 6]). From the structural perspective, many important graph classes such as hyperbolic graphs [20], outerstring graphs [23, 4], and (theta,prism,pyramid)-free graphs [1, 12] have bounded strong isometric path complexity [6]. It was shown in [6, 8] that $K_{2,3}$-induced minor-free graphs have bounded strong isometric path complexity. We shall use this fact in the proof of Theorem 3. On the other hand, graphs with tree-width at most two have unbounded strong isometric path complexity.
@@
Overview of the proof of Theorem 3: Given a $K_{2,3}$-induced minor free graph $G$, we choose a vertex $r\in V(G)$ arbitrarily as the root and construct the layering tree w.r.t $r$, denoted by $\Gamma\left({r},{G}\right)$. Let $R=\{r\}$ be the root cluster. When $\Gamma\left({r},{G}\right)$ is rooted at $R$, there is a natural parent-child relationship among the clusters. For a cluster $S$, the parent set of $S$, denoted as $P\left(r,S\right)$, is the set of vertices that lie in the parent of $S$ and are adjacent to the vertices in $S$. We observe that given a cluster $S$ whose parent is not $R$, $P\left(r,S\right)$ is a minimal cutset of the original input graph. The size of the largest stable set in the graph induced by a minimal cutset of $K_{2,3}$-induced minor-free graph is at most two [15]. Hence, for any cluster $S$ whose parent is not $R$, the size of the largest stable set in the graph induced by $P\left(r,S\right)$ in $G$ is at most two.
@@
We begin our construction by initializing a graph $H=(V(G),\emptyset)$. We traverse $\Gamma\left({r},{G}\right)$ (rooted at $R$) bottom-up and process the clusters in $G$ accordingly. For each cluster $S$, depending on the structure of the graph induced by $P\left(r,S\right)$, we introduce edges in $H$ between the vertices of $S$ and at most two special vertices of $P\left(r,S\right)$. See Algorithm 1 for details of the construction of $H$. Note that the final resulting graph $H$ may not be a subgraph of $G$.
@@
To prove the upper bound on the tree-width of $H$, we show that $H$ is triangle-free and does not contain wheel or theta as induced subgraphs. See Section 2.1 for definitions. The absence of these structures in $H$ implies that $H$ is a universally signable graph [13], and has tree-width at most two. This concludes the proof of Theorem 3.
Organisation: In Section 2, we introduce some notations, recall some existing results, and prove some preliminary observations. In Section 3, we prove Theorem 3 and conclude in Section 4.
@@ 2 Preliminaries and some observations
All graphs in this paper are finite, undirected, unweighted, and simple. Let $G$ be a graph. For a set $S\subseteq V(G)$, $G\left[S\right]$ denotes the subgraph of $G$ induced by $S$, and $G-S=G\left[V(G)\setminus S\right]$. When $S=\{w\}$, we write $G-\{w\}$ as $G-w$. For two graphs $G,H$, we write $H\subseteq G$, if $H$ is a subgraph of $G$. A stable set of a graph is a set of pairwise non-adjacent vertices. A cutset of a graph $G$ is a set of vertices of $G$ whose removal disconnects $G$. A cutset $S$ of a graph $G$ is minimal if no proper subset of $S$ is a cutset of $G$.
@@ 2.1 $K_{2,3}$-induced minor-free graphs
The size of the largest stable set of the graph induced by any minimal cutset of a $K_{2,3}$-induced minor-free graph is at most two.
@@
Universally signable graphs form a subclass of $K_{2,3}$-induced minor-free graphs. See [13] for the original definition of universally signable graphs. A graph is universally signable if and only if it does not contain any theta, pyramid, prism, or wheel [13]. In the same paper, it was also proved that any minimal cutset in a universally signable graph G is either a clique or a stable set of size two. The next proposition follows from the proof of the above result.
@@
Proposition 3 implies the following.
@@ 2.2 Paths, distances and layering partition
Below, we show that the parent set of every cluster is a minimal cutset except in one case.
@@
Let $G$ be a graph, $r$ be a vertex, and $S$ be a cluster w.r.t $r$ such that $P\left(r,S\right)\neq\{r\}$. Then $P\left(r,S\right)$ is a minimal cutset of $G$.
@@
Let $X\coloneq P\left(r,S\right)$ and $S\subseteq N_{k}(r)$ with $k\geq 2$. For any $u\in S$, an $(r,s)$-path in $G$ must contain a vertex of $X$, and therefore, $r$ and $u$ lie in different connected components in $G-X$. Hence, $X$ is a cutset.
Let $Y\subset X$, $H=G-Y$, $u\in X\setminus Y$, and $v\in S$ with $uv\in E(H)$. Let $Z_{1}$ be the set of vertices that lie in $S$ or in some cluster which is a descendent of $S$ in $\Gamma\left({r},{G}\right)$. Observe that $Z_{1}\subset V(H)$ and in particular, $v\in Z_{1}$. Since any vertex $w\in Z_{1}$ is connected to $v$ by a path in $H\left[Z_{1}\right]$, we have that $H\left[Z_{1}\right]$ is connected. Let $Z_{2}=V(G-Z_{1}-Y)$. By definition, $V(H)=Z_{1}\cup Z_{2},r\in Z_{2}$ and $u\in Z_{2}$. Observe that for any vertex $w\in Z_{2}$, there is a $(r,w)$-path in $H$. Hence $H\left[Z_{2}\right]$ is connected. As $uv\in E(H)$, $H$ is connected. ∎
Lemma 4 and Proposition 1 imply the following properties of clusters of $K_{2,3}$-induced minor-free graphs.
@@
Let $G$ be a $K_{2,3}$-induced minor-free graphs, $r$ be a vertex, $S$ be a cluster w.r.t $r$, and $F\coloneq G\left[P\left(r,S\right)\right]$. Then all of the following holds. (a) $F$has at most two connected components, (b) If $F$ is connected then it has diameter at most three, (c) If $F$ has two connected components, then each connected component is a clique.
@@
If there exists a connected component $Y\in\{C_{1},D_{1}\}$ containing vertices from both $C_{2}$ and $D_{2}$, then due to Lemma 5(c), there will be an edge in $G\left[P\left(r,S_{2}\right)\right]$ between vertices of $C_{2}$ and $D_{2}$, which is a contradiction. Now assume $|\mathcal{X}|\geq 2$. Without loss of generality, let $C_{1}$ contains a vertex $c$ from $C_{2}$. Then, $D_{1}$ must contain a vertex $d$ from $D_{2}$. For $i\in\{1,2\}$, let $c_{i}\in N(c)\cap S_{i}$ and $d_{i}\in N(d)\cap S_{i}$. Let $k=d\left(r,S\right)$. For $i\in\{1,2\}$, let $P_{i}$ denote a $(c_{i},d_{i})$-induced path in $G-B_{r}\left({k}\right)$. (Since $c_{i},d_{i}$ lie in the same cluster in $G$, $P_{i}$ exists). Let $P$ be an $(c,d)$-upper path w.r.t $r$. the graph induced by the vertices of $P,P_{1},P_{2}$, contains a theta, which contradicts Proposition 2. ∎

### v2

@@ Abstract
A graph $H$ is an induced minor of a graph $G$ if $H$ can be obtained from $G$ by a sequence of edge contractions and vertex deletions. Otherwise, $G$ is $H$-induced minor-free. In this paper, we provide a different proof of the fact that $K_{2,3}$-induced minor-free graphs admit a quasi-isometry with additive distortion to graphs with tree-width at most two. Our proof yields a $O(nm)$-time algorithm which takes as input a $K_{2,3}$-induced minor-free graph with $n$ vertices and $m$ edges, and outputs a tree-width two graph $H$ with the desired additive distortion. For universally signable graphs, a subclass of $K_{2,3}$-induced minor-free graphs, the time complexity of our algorithm is linear. As a consequence, we obtain a truly sub-quadratic time additive constant factor approximation algorithm to compute the diameter of a universally signable graph. In contrast, assuming the Strong Exponential Time Hypothesis (SETH), the diameter of split graphs (a very restricted class of universally signable graphs), cannot be computed in truly sub-quadratic time [Borassi et al. (ENTCS, 2016)].
@@ 1 Introduction
When is a graph quasi-isometric (Definition˜1) to a much simpler host graph? This question is central to coarse graph theory [22]. When the host graphs are restricted to the class of graphs with bounded tree-width, remarkably, the answer is provided by Nguyen, Scott & Seymour [26] and independently by Hickingbotham [24]. Specifically, they proved the following:
@@
When quasi-isometric maps with only additive distortions (Definition˜2) are allowed, the situation is much less understood. Brandstädt et al. [6] proved that chordal graphs (i.e., graphs without any induced cycle of length greater than three) admit a quasi-isometry to trees with additive distortion at most $2$. This result has been later generalized to $k$-chordal graphs [12], and more generally to graphs of bounded treelength [18]. Berger & Seymour [3] characterized graphs admitting quasi-isometry with additive distortion to trees. See [19] for a comprehensive overview of graphs admitting a quasi-isometry with additive distortion to trees.
@@
In this paper, we consider the class of $K_{2,3}$-induced minor-free graphs. A graph $H$ is an induced minor of a graph $G$ if $H$ can be obtained from $G$ by a sequence of edge contractions and vertex deletions. Otherwise, $G$ is $H$-induced minor-free. The class of $K_{2,3}$-induced minor-free graphs have been studied extensively in structural graph theory [16, 17]. $K_{2,3}$-induced minor-free graphs generalize many popular graph classes. Examples include outerplanar graphs, circular-arc graphs [20], chordal graphs, universally signable graph [15] etc. Coarse generalizations of $K_{2,3}$-induced minor-free graphs have been studied [13]. A recent result of Albrechtsen et al. [2] implies that $K_{2,3}$-induced minor-free graphs are quasi-isometric to graphs with tree-width at most two. Combining results of Fujiwara & Papasoglu [21] and Nguyen, Seymour & Scott [26] prove that ˜2 holds for $K_{2,3}$-induced minor-free graphs [^2] We thank Robert Hickingbotham for informing us of this connection. In this paper, we provide a different constructive proof for the above. Specifically, we prove the following theorem.
@@
There is a constant $c$, and an $O(nm)$-time algorithm that takes as input a $K_{2,3}$-induced minor-free graph $G$ with $n$ vertices and $m$ edges, and outputs a graph $H$ with $V(G)=V(H)$, tree-width at most two, and $\left|d_{G}\left(u,v\right)-d_{H}\left(u,v\right)\right|\leq c$ for any $u,v\in V(G)$.
In the above theorem, $d_{G}\left(.,.\right)$ and $d_{H}\left(.,.\right)$ denote the standard graph metric in $G$ and $H$, respectively. When the input graphs are restricted to universally signable graphs [15], a subclass of $K_{2,3}$-induced minor-free graph, our algorithm runs in linear time, which has the following consequence.
The diameter of a graph is the length of a longest isometric path (i.e. a path of minimum length between the end-vertices) between two of its vertices. Given an undirected graph $G$, Diameter is the algorithmic task of computing the diameter of $G$. Diameter is a fundamental problem with countless applications in computer science. The fastest combinatorial algorithm for Diameter has a running time of $O(nm)$ in $m$-edge, $n$-vertex graphs. Surprisingly, Roditty and Vassilevska Williams [27] proved that a truly sub-quadratic $\left(\frac{3}{2}-\epsilon\right)$-approximation algorithm for Diameter would falsify the Strong Exponential Time Hypothesis [25]. This motivated researchers to study the approximability of Diameter on special graph classes. However, Borassi, Crescenzi & Habib [4] proved that assuming SETH, diameter cannot be computed optimally in truly sub-quadratic time even on split graphs, i.e. graphs whose vertex set can be partitioned into a clique and an independent set. Split graphs are universally signable graphs. Hence assuming SETH, the diameter of universally signable graphs cannot be computed optimally in truly sub-quadratic time. The diameter of graphs with tree-width at most two can be computed in truly sub-quadratic time [7]. Hence, Theorem˜3 imply the following corollary.
Corollary 1.
Diameter admits a truly sub-quadratic additive constant factor approximation algorithm on universally signable graphs.
@@
Below, we provide an informal description of the tools used in the proofs of Theorem˜3. Then, we provide an overview of the proofs.
@@
Important tools: The proof of Theorem˜3 uses layering partition [6, 12] and the notion of strong isometric path complexity [8] of graphs. Layering partitions have been used to solve numerous problems in metric graph theory [6, 12, 13, 19]. See Section˜2.2 for definitions. Informally, in a layering partition, the vertex set of a graph is partitioned into clusters w.r.t a fixed root vertex. Two vertices are in the same cluster if they are at a distance $k$ from the root for some $k\geq 0$ and are end-vertices of a path that is disjoint from the ball of radius k1k-1 centered at the root vertex. It was shown in [6, 12] that the adjacency relation between the clusters is captured by a tree, known as the layering tree. Moreover, for a fixed root vertex, the corresponding layering tree can be constructed in linear time [12].
@@
The strong isometric path complexity is a recently introduced graph invariant that captures how arbitrary isometric paths of a graph can be viewed as a union of few “rooted" isometric paths (i.e. isometric paths with a common end-vertex). See Definition˜4. Informally, if the strong isometric path complexity of a graph $G$ is at most $k$, then for any isometric path $P$ of $G$ and any vertex of $v\in V(G)$, the vertices of $P$ can be covered with $k$-many $v$-rooted isometric paths of $G$. Originally, strong isometric path complexity and its weaker variant isometric path complexity were introduced to propose approximation algorithms for the problem Isometric Path Cover, where the objective is to find a minimum-cardinality set of isometric paths so that each vertex of the graph belongs to at least one of the solution paths. Isometric Path Cover has been studied recently in the algorithmic graph theory community [9, 8, 11]. It was shown that Isometric Path Cover admit a polynomial-time $c$-approximation algorithm on graphs with strong isometric path complexity at most $c$ [8] (while being NP-hard even on chordal graphs whose strong isometric path complexity is at most four [9, 8]). From the structural perspective, many important graph classes such as hyperbolic graphs [23], outerstring graphs [28, 5], and (theta,prism,pyramid)-free graphs [1, 14] have bounded strong isometric path complexity [8]. It was shown in [8, 10] that $K_{2,3}$-induced minor-free graphs have bounded strong isometric path complexity. We shall use this fact in the proof of Theorem˜3. On the other hand, graphs with tree-width at most two have unbounded strong isometric path complexity.
@@
Overview of the proofs of Theorem˜3: Given a $K_{2,3}$-induced minor free graph $G$, we choose a vertex $r\in V(G)$ arbitrarily as the root and construct the layering tree w.r.t $r$, denoted by $\Gamma\left({r},{G}\right)$. Let $R=\{r\}$ be the root cluster. When $\Gamma\left({r},{G}\right)$ is rooted at $R$, there is a natural parent-child relationship among the clusters. For a cluster $S$, the parent set of $S$, denoted as $P\left(r,S\right)$, is the set of vertices that lie in the parent of $S$ and are adjacent to the vertices in $S$. We observe that given a cluster $S$ whose parent is not $R$, $P\left(r,S\right)$ is a minimal cutset of the original input graph. The size of the largest stable set in the graph induced by a minimal cutset of $K_{2,3}$-induced minor-free graph is at most two [17]. Hence, for any cluster $S$ whose parent is not $R$, the size of the largest stable set in the graph induced by $P\left(r,S\right)$ in $G$ is at most two.
@@
We begin our construction by initializing a graph $H=(V(G),\emptyset)$. We traverse $\Gamma\left({r},{G}\right)$ (rooted at $R$) bottom-up and process the clusters in $G$ accordingly. For each cluster $S$, depending on the structure of the graph induced by $P\left(r,S\right)$, we introduce edges in $H$ between the vertices of $S$ and at most two special vertices of $P\left(r,S\right)$. See Algorithm˜1 for details of the construction of $H$. Note that the final resulting graph $H$ may not be a subgraph of $G$.
@@
To prove the upper bound on the tree-width of $H$, we show that $H$ is triangle-free and does not contain wheel or theta as induced subgraphs. See Section˜2.1 for definitions. The absence of these structures in $H$ implies that $H$ is a universally signable graph, and has tree-width at most two. This concludes the proof of Theorem˜3.
Organisation: In Section˜2, we introduce some notations, recall some existing results, and prove some preliminary observations. In Section˜3, we prove Theorem˜3 and conclude in Section˜4.
@@ 2 Preliminaries and some observations
All graphs in this paper are finite, undirected, unweighted, and simple. Let $G$ be a graph. For a set $S\subseteq V(G)$, $G\left[S\right]$ denotes the subgraph of $G$ induced by $S$, and $G-S=G\left[V(G)\setminus S\right]$. When $S=\{w\}$, we write $G-\{w\}$ as GwG-w. For two graphs $G,H$, we write $H\subseteq G$, if $H$ is a subgraph of $G$. A stable set of a graph is a set of pairwise non-adjacent vertices. A cutset of a graph $G$ is a set of vertices of $G$ whose removal disconnects $G$. For two vertices $u,v\in V(G)$, a $(u,v)$-cutset of a graph $G$ is a set of vertices of $G$ whose removal disconnects $u,v$ in $G$. A cutset $S$ of a graph $G$ is minimal if no proper subset of $S$ is a cutset of $G$. Similarly, for two vertices $u,v\in V(G)$, a $(u,v)$-cutset $S$ of a graph $G$ is minimal if no proper subset of $S$ is a $(u,v)$-cutset of $G$.
@@ 2.1 $K_{2,3}$-induced minor-free graphs
Let $G$ be a $K_{2,3}$-induced minor-free graph. For any two vertices $u,v\in V(G)$, the size of the largest stable set of the graph induced by any minimal $(u,v)$-cutset of $G$ is at most two.
@@
Universally signable graphs form a subclass of $K_{2,3}$-induced minor-free graphs. See [15] for the original definition of universally signable graphs. A graph is universally signable if and only if it does not contain any theta, pyramid, prism, or wheel [15]. Important graph classes like chordal graphs, outerplanar graphs are sub classes of universally signable graphs. Any minimal cutset in a universally signable graph G is either a clique or a stable set of size two [15]. The next proposition follows from the proof of the above result.
@@
Proposition˜3 implies the following.
Proposition 4.
Let $G$ be a universally signable graph and $C$ be an induced cycle of $G$ with length greater than three. Let $u,v\in C$ be two non-adjacent vertices of $P$. Then for any indcued path $P$ between $u$ and $v$, we have $V(P)\subseteq V(C)$.
@@ 2.2 Paths, distances and layering partition
Below, we show that the parent set of every cluster is a minimal $(r,v)$-cutset where $v$ is a vertex of the cluster.
@@
Let $G$ be a graph, $r$ be a vertex, $S$ be a cluster w.r.t $r$ and $u\in S$ such that $P\left(r,S\right)\neq\{r\}$. Then $P\left(r,S\right)$ is a minimal $(r,u)$-cutset of $G$.
@@
Let $X\coloneq P\left(r,S\right)$ and $S\subseteq N_{k}(r)$ with $k\geq 2$. For any $u\in S$, an $(r,s)$-path in $G$ must contain a vertex of $X$, and therefore, $r$ and $u$ lie in different connected components in GXG-X. Hence, $X$ is a cutset.
Let $Y\subset X$, H=GYH=G-Y, $u\in X\setminus Y$, and $v\in S$ with $uv\in E(H)$. Let $Z_{1}$ be the set of vertices that lie in $S$ or in some cluster which is a descendent of $S$ in $\Gamma\left({r},{G}\right)$. Observe that $Z_{1}\subset V(H)$ and in particular, $v\in Z_{1}$. Since any vertex $w\in Z_{1}$ is connected to $v$ by a path in $H\left[Z_{1}\right]$, we have that $H\left[Z_{1}\right]$ is connected. Let $Z_{2}=V(G-Z_{1}-Y)$. By definition, $V(H)=Z_{1}\cup Z_{2},r\in Z_{2}$ and $u\in Z_{2}$. Observe that for any vertex $w\in Z_{2}$, there is a $(r,w)$-path in $H$. Hence $H\left[Z_{2}\right]$ is connected. As $uv\in E(H)$, $H$ is connected. ∎
Lemma˜4 and Proposition˜1 imply the following properties of clusters of $K_{2,3}$-induced minor-free graphs.
@@
Let $G$ be a $K_{2,3}$-induced minor-free graphs, $r$ be a vertex, $S$ be a cluster w.r.t $r$, and $F\coloneq G\left[P\left(r,S\right)\right]$. Then all of the following holds. (a) FFhas at most two connected components, (b) If $F$ is connected then it has diameter at most three, (c) If $F$ has two connected components, then each connected component is a clique.
@@
If there exists a connected component $Y\in\{C_{1},D_{1}\}$ containing vertices from both $C_{2}$ and $D_{2}$, then due to Lemma˜5(c), there will be an edge in $G\left[P\left(r,S_{2}\right)\right]$ between vertices of $C_{2}$ and $D_{2}$, which is a contradiction. Now assume $|\mathcal{X}|\geq 2$. Without loss of generality, let $C_{1}$ contains a vertex $c$ from $C_{2}$. Then, $D_{1}$ must contain a vertex $d$ from $D_{2}$. For $i\in\{1,2\}$, let $c_{i}\in N(c)\cap S_{i}$ and $d_{i}\in N(d)\cap S_{i}$. Let $k=d\left(r,S\right)$. For $i\in\{1,2\}$, let $P_{i}$ denote a $(c_{i},d_{i})$-induced path in $G-B_{r}\left({k}\right)$. (Since $c_{i},d_{i}$ lie in the same cluster in $G$, $P_{i}$ exists). Let $P$ be an $(c,d)$-upper path w.r.t $r$. the graph induced by the vertices of $P,P_{1},P_{2}$, contains a theta, which contradicts Proposition˜2. ∎